### v1

# Long-Tail Learning with Rebalanced Contrastive Loss

Abstract
Integrating supervised contrastive loss to cross entropy-based classification has recently been proposed as a solution to address the long-tail learning problem. However, when the class imbalance ratio is high, it requires adjusting the supervised contrastive loss to support the tail classes, as the conventional contrastive learning is biased towards head classes by default. To this end, we present Rebalanced Contrastive Learning (RCL), an efficient means to increase the long tail classification accuracy by addressing three main aspects: 1. Feature space balancedness – Equal division of the feature space among all the classes 2. Intra-Class compactness – Reducing the distance between same-class embeddings 3. Regularization – Enforcing larger margins for tail classes to reduce overfitting. RCL adopts class frequency-based SoftMax loss balancing to supervised contrastive learning loss and exploits scalar multiplied features fed to the contrastive learning loss to enforce compactness. We implement RCL on the Balanced Contrastive Learning (BCL) Framework, which has the SOTA performance. Our experiments on three benchmark datasets demonstrate the richness of the learnt embeddings and increased top-1 balanced accuracy RCL provides to the BCL framework. We further demonstrate that the performance of RCL as a standalone loss also achieves state-of-the-art level accuracy.
Keywords:
Long-tail Classification Contrastive Learning1 Introduction
The data imbalance between different classes usually hinders training an unbiased classifier. This is a significant problem, especially when the tail class is of great importance for applications such as pedestrian classification, fraud detection, and spam filtering. While approaches such as data resampling and loss reweighting are used to alleviate class imbalance, these approaches, however, increase the accuracy of the tail classes by sacrificing the head class accuracy. An important aspect of a robust model is the capability to address the distribution shift from training (long-tail) to unknown test (uniform or inversely long-tailed) distributions. Later in some frameworks, cross-entropy loss for long-tail learning is combined with metric learning or contrastive learning. Especially work such as [12] demonstrates strong performance improvement when the cross-entropy loss is combined with Supervised Contrastive Learning (SCL). This work demonstrates that learning better embeddings can assist in resolving most of the long-tail classification challenges, as better embeddings can contribute to more generalized classifiers. Along this line, one approach is to decouple representation learning and classifier learning [14]. However, SCL [15] based learnt representation can not eliminate domination by most frequent classes and the decrease of linear separability of the learnt embeddings from head to tail classes unless necessary adjustments for long-tail distributions are conducted. Therefore, it is important to improve SCL to handle long-tail data distributions more robustly.
Zhu et al. [29] propose balanced contrastive loss BCL, an improved SCL loss for the context of long-tail learning where the main intuition is to average the attraction and repulsion terms based on the mini-batch class frequency so that each class equally contributes to the optimization process. We propose a means to further enhance the quality of the learnt embedding of BCL from the perspectives of balancedness, intra-class compactness, and regularization using Rebalanced Contrastive Learning (RCL). We emphasize that the RCL is not limited to working with BCL only. Given its simplicity of implementation, it can be incorporated into any SCL method. More specifically, we devise the following contributions.
We propose RCL to optimize the learnt representation of BCL through i) establishing the robustness for distribution shift from train (long tail) to test (uniform) using the adopted class frequency-based SoftMax balancing, ii) enforcing intra-class compactness for tail classes through scalar multiplied features fed to the contrastive loss, and iii) enforcing better class separation through enforced class frequency based embedding margins.
By conducting experiments on the long-tail versions of the three benchmark datasets: CIFAR10, CIFAR100, and ImageNet, we show that the addition of RCL over BCL improves the top-1 balanced accuracy. Our ablation study further shows that RCL as a standalone metric also increases the top-1 harmonic mean accuracy, indicating improvements in the least-performing classes.
We conduct qualitative evaluations of the learnt embeddings and demonstrate that adding RCL to BCL enhances class separability with larger margins through balancedness, intra-class compactness, and regularization.
2 Related Work
We explain related work in long-tail classification under three topics: i) Data-level solutions, ii) Algorithm-level solutions, and iii) SCL-based approaches.
Data-level solutions - Early works in long-tail learning include data-level solutions such as random oversampling, random undersampling [3], square root sampling [19] and progressively balanced sampling [14] that suffer from over-fitting or under-fitting to data. Re-weighting methods [8, 25] assign weights on different training samples based on the label or instance (hard or easy samples) and have shown significant accuracy increments over few-shot classes. However, re-weighting methods tend to sacrifice the many-shot accuracy. Classifier balancing techniques such as fine-tuning the learnt classifier [10], post-hoc calibrations [20, 14], domainshift calibrations [21] alleviate the imbalances of the data. However, these methods may not be sufficient under extreme class imbalance ratios and when the number of classes is very high. Data augmentation is also used to attain significant accuracy improvements for classes suffering from absolute rarity [22]. Works such as [5] use generative models to generate new samples for tail classes as a convex combination of existing samples. Transferring the information from head classes to tail classes also provides promising results [16]. Yet, these methods can not completely eliminate the tail class over-fitting and the dominance of head classes.
Algorithmic-level solutions - Under algorithmic level solutions integrating a per-class margin [20, 23] into the cross-entropy loss is a popular direction of work. In most cases, the logit margin is set as inversely proportional to the class frequency to enforce a larger margin for tail classes to enhance generalization. However, these methods do not control the learned feature distributions explicitly. Therefore, it is important to use techniques to optimize the learned feature distribution to avoid the diffusion of the rare classes.
Representation optimization with SCL - The contrastive learning-based methods have been used recently in the context of long-tail learning as they contribute to better feature representation by aligning the same class pairs together and repulsing the different class pairs. In [18], authors propose a novel technique to improve SCL so that it computes optimal class centres to balance the features space evenly. Then, they associate the generated class centres with actual class IDs considering mutual closeness. However, the results here could be suboptimal since there is no deterministic strategy to associate the class centres with the true classes. Song et al. [26] propose an improved SCL framework denoted as prototypical supervised contrastive (PSC) learning where it learns prototypes for each class; subsequently, each instance is to be pushed towards the prototypes of its corresponding class and pulled away from the other class prototypes. In this approach, tail class instances may have a larger effect on the model optimization compared to head class instances, sacrificing the head class accuracy. To address this issue, [11] proposes subclass-balancing contrastive learning where the loss is defined on subclasses equal to the tail class size inside the head classes through clustering. Here, the authors address issues with balancing the SCL loss minimized along with the CE loss in the context of long-tail classification scenarios with higher class imbalance. They focus on increasing the balancedness, intra-class compactness, and tail-class margins through rebalancing the BCL loss.
3 Methodology
3.1 Notations
Multiclass classification problem can be defined by input instance domain , class label domain is given by . For a set of training samples where joint distribution P characterized on . All the weights corresponding to the model are denoted by . The multi-class learning task is to learn a scorer so that it minimizes the expected loss . The scorer corresponding to class y is , where are the weights of the last fully connected classification layer corresponding to class y. And is the bias. Learned embedding corresponding to the input is . Further, we denote the batch frequency of class y by and the learnt class prototype by as defined in [29].
3.2 Preliminaries
CE Based Classification: Our work makes use of fundamental Cross Entropy (CE) modification methods for long-tail learning. Firstly, as the SoftMax function is inherently biased under long-tail learning, a Balanced SoftMax function [23] is proposed to address the train-to-test distribution shift from a probabilistic perspective. The modified loss is denoted in Equation 1. The authors provided an analytical derivation that optimizing for the Balanced SoftMax cross-entropy loss is essentially equivalent to minimizing the bound on generalization error. The second interesting approach is logit adjustment (LA) proposed by Menon et al. [20] that allows larger margins for tail classes, which can be denoted with Equation 2. Here denotes the probability of class y. This model is more favourable as provides the degree of freedom to adjust for different datasets and class imbalance levels.
SCL Integration: Above LA cross entropy classification module combined with a supervised contrastive learning module to enhance the long-tail classification accuracy. The supervised contrastive learning component is balanced through class averaging and class complements [29]. The new loss is denoted as balanced contrastive loss BCL. This loss avoids the bias occurred towards the head classes. Class complements is an effort to have all classes to be present in every mini batch through the use of class centre prototypes. When the rare class instances are missing from a batch the centre embedding would act as an actual training instance. Refer to the BCL loss component given in Equation 3 for the use of class averaging and class complements.
Our proposed approach RCL is implemented on BCL to incorporate more balancedness through the adopted balanced SoftMax concept and enforce larger margins for tail classes. The BCL + RCL is integrated with the logit-adjusted CE loss to conduct the classification. Furthermore, RCL loss can be implemented on different supervised contrastive losses along with any combination of cross-entropy loss with class imbalance adjustment.
3.3 Feature Representation Optimization with RCL
Our RCL-based approach addresses three key aspects in the embedding space that can directly contribute to increasing classification accuracy: 1) Balanced feature space, 2) Intra-class compactness, and 3) Assigning larger margins to ambiguous classes. These key aspects are expected to provide more support for tail class classification. Figure 0(a) denotes the distribution of the learned embedding under the class imbalance for a simple separable 2D problem. The majority class tend to occupy the larger area of the feature space with larger margins, thus allowing larger generalisation on the head classes and over-fitting to minority classes. In Figure 2, we illustrate the pipeline for the RCL-based learning. In this work, we apply the RCL loss to the BCL loss. However, RCL loss can be integrated with any given similar SCL pipeline.
3.3.1 Balancing Feature Space
It was noted that unsupervised contrastive learning methods are more robust to highly imbalanced class distributions compared to supervised contrastive learning [13]. The reason for this phenomenon is that the unsupervised CL methods generate a balanced feature space for all the classes compared to SCL. To this end, some research focuses on pushing the learnt embeddings into pre-defined class centres to obtain the balance [18]. In this work, we try to obtain balancedness by balancing the SCL loss sufficiently. The modified SCL loss is defined as Rebalanced Contrastive Loss (RCL). We adopt the Balanced SoftMax concept described in Section 3.2 to modify the SCL loss as indicated in Equation 4. Where is the frequency of the class y in the complete training set. The RCL modification can also be implemented on BCL to enhance the balancedness as in Equation 5. The fraction denoted in red is the class averaged similarity scores.
3.3.2 Enforcing feature compactness
Minimizing the intra-class distance among the same class embeddings is a main aspect of interest as it reduces the deviation of the features across the decision boundary and enhances class discriminativeness. Feature Clusters Compression (FCC) [17] is applied previously to decrease the diffusion of backbone features by compressing learnt embeddings. This is simply achieved by multiplying original learnt features by a selected scaling factor and using the multiplied features in computing the cross entropy loss for classifier learning. In this work, we modify this approach and integrate the feature compression component directly into the RCL instead of sending it through the classifier as in CE minimization, i.e., . The features are scalar multiplied before feeding to the SCL loss. The intuition is to provide a higher degree of freedom to reshape features based on feature compression. The scaling factor and value is selected based on the class id. At the The validation accuracy is obtained, and if the validation per class accuracy is lower than , a novel scaling factor is applied for the corresponding classes for the rest of the classes, the factor is set to 1 (only two different values are used for the entire data set). The scaling factor used on the underperforming classes is selected based on a grid search.
3.3.3 Regularization
The rare classes must be more strongly regularized compared to frequent classes to reduce the generalization error of the rare classes without sacrificing the accuracy of majority classes. To achieve this it requires a data-dependant or label-dependent regularizer rather than the standard L2 regularize, which only depends on the model weights. Therefore, the data-dependent properties like class margins can be useful. Enforcing a larger margin for a class can be interpreted as regularization [2] because the inverse of the minimum margin among all the examples is linked with the standard generalization error bounds. According to [12], enforcing logit margins is insufficient under extreme imbalance scenarios, and therefore, embedding margins are also used to enhance the regularization of the minority classes. We demonstrate the importance of the RCL in the context of enforced embedding margins by reformulating it as in Equation 6 followed by Equation 7. According to this format, it is evident that margin is implicitly enforced for the similarity score between the same class and dissimilar class embeddings based on the class frequency. We can further analyse the loss to get a comprehensive insight considering the pairwise margin loss.
Pairwise margins denoted in Equation 7 enforces the desired gap between the similarity scores and depending on the belonginess to same or different classes. This imposes a larger margin between the tail intra-class similarity score and tail and head inter-class similarity score so that it enhances the feature compactness within the tail class and increases the distance between head and tail feature clusters. Figure 0(b) denotes a 2D classification problem which portrays the ideally expected distribution of the learnt features under an imbalanced data distribution. All the classes occupy a similar area in the feature space, and the features are well compressed towards the class centre while providing larger margins for the rare classes.
4 Experiments
| LC | SCL | BCL | RCL | Top-1 Accuracy (Arithmetic) | Top-1 Accuracy (Harmonic) |
[…]
| Dataset | CIFAR10 | CIFAR100 | ImageNet |
| Imbalance Ratio | 100 | 100 | 256 |
| CE | 70.4 | 38.5 | 41.6 |
| LDAM+DRW [2] | 77.2 | 42.2 | 49.8 |
| LogitAdjust [20] | 80.2 | 43.9 | 50.9 |
| CB+Focal [8] | 74.6 | 39.9 | - |
| BALMS [23] | 84.9 | 50.8 | - |
| Two Stage Methods | |||
| Decoupled [14] | 78.5 | 41.3 | 50.6 |
| DRO_LT [24] | 80.5 | 47.1 | 53.5 |
| Ensemble Methods | |||
| RIDE [28] | - | 49.1 | 55.4 |
| ACE [1] | 81.4 | 49.6 | 54.7 |
| CE+ other Loss | |||
| ELM [12] | 78.0 | 45.8 | 50.6 |
| CE+ Supervised Contrastive Loss | |||
| Hybrid_spc [27] | 81.4 | 46.7 | - |
| BCL | 84.3 | 51.9 | 56.2 |
| RCL (ours) | 86.0 | 52.2 | 55.4 |
| BCL+RCL (ours) | 86.2 | 52.1 | 56.3 |
To validate our findings on representation learning optimization with RCL, we use three benchmark datasets containing significant class imbalance, exhibiting long-tail distribution and involving a large number of classes. We next describe the datasets, experiment design, and performance metrics, followed by comparisons of the results with the SOTA techniques and ablation studies.
4.1 Datasets
We use the long-tail versions of the benchmark datasets: CIFAR10, CIFAR100, and ImageNet. The level of class imbalance is measured by the Class Imbalance Factor (IF), which is the ratio between the most frequent class frequency and the least frequent class frequency. We set the maximum possible imbalance factors for the datasets, leaving at least 5-6 instances per tail class. Doing so, we obtain the following experiment settings; i) CIFAR10: No. of classes 10, IF - 100, majority class frequency - 5,000, ii) CIFAR100: No. of classes 100, IF - 100, majority class frequency - 500, iii) ImageNet: No. of classes 1,000, IF - 256, majority class frequency - 1,280.
4.2 Implementation
Referring to the previous implementation, we use ResNet32 backbone for the CIFAR datasets and ResNext-50 and ResNet50 for the ImageNet_Lt. Models are trained with a batch size of 256. Learning rate for CIFAR10 and 100 are set to 0.2 and for ImageNet 0.1. Similar to [7] and [29], we use Autoaugment [6] and Cutout [9] for augmenting the features fed to the classifier. SimAugment [4] is used for the features fed to the contrastive loss. The model training is conducted on the Amazon Cloud with RONIN mediation service with 8xV100 GPUs.
4.3 Ablation Study
We first conduct an ablation study on CIFAR100 with IF set to 100 to analyse the performance benefits of RCL. The overall balanced accuracy (averaged per class) is reported in Table 1. As small values heavily influence arithmetic means in obtaining the overall class accuracy, we also report the harmonic mean to reflect the performance of the worst-performing classes. Initially, the dataset is tested for cross-entropy loss with LA [20]. Where the logit adjustment parameter is set to 1.3 through grid search. Then, the accuracy is evaluated by combining standard supervised contrastive loss with setting to 2 and to 1 and using SimAugment for augmentation. The compression factor is set to 0.005 for the classes with validation accuracy less than , and for the other classes, the factor is set to 1. Compression is applied from the epoch.
Results in Table 1 show adding RCL to standard SCL would increase top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4). When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5). SCL+RCL demonstrates the highest harmonic mean accuracy of all the combinations.
4.4 Performance Analysis
4.4.1 Baselines
We compare our proposed framework with twelve SOTA models. Balanced overall test accuracy (arithmetic mean over per-class accuracy) is used for comparison. The selected baselines are: 1) standard cross entropy minimisation, 2) LDAM-DRW where logit margin is enforced depending on the class frequency and reweighting scheme is used in optimization [2], 3) Class balanced focal loss [8] which assigns a frequency based weighting on per class losses, 4) Additive logit margin based on the class frequency [20], 5) BALMS - Balanced SoftMax function Equation [23] to address the distribution shift integrated with a meta sampler that can compute the optimal class sample rate, 6) Decoupled learning which enhances the classification accuracy through fine-tuning a learnt classifier with class-balanced sampling or classifier weight normalization [14], 7) robustness theory based loss to enhance the classification and feature representation [24], 8) RIDE [28] and 9) ACE [1] which are multi expert ensemble models, 10) ELM [12] which enforces logit and embedding margins, and 11) Hybrid_spc [27] prototypical supervised contrastive (PSC) learning.
4.4.2 Results
We show the results in Table 2. As can be seen, standalone RCL demonstrates , improvement over CIFAR10 and 100, respectively, compared to BCL, which is the next best model. Furthermore, BCL+RCL has consistent improvements , , and over all the datasets in comparison to BCL. In summary, standalone RCL and BCL+RCL outperform all the other baselines apart from one scenario where BCL performs best (i.e., ImageNet).
4.4.3 Embedding space analysis
In Figure 3, we show the TSNE plots of the learnt representation of the model for CIFAR10 and ImageNet. BCL+RCL demonstrates a better class separation and balanced division of the feature space compared to BCL. As seen from Figure 3(a), in BCL, tail class embeddings are grouped into sub-clusters and lie far apart. Once the features compression is added, all embeddings corresponding to the same tail class are grouped together, providing better intra-class compactness and balancedness in the space in Figure 3(b). For example, in Figure 3(a), tail classes 8 and 9, represented by orange and cyan, have two sub-clusters each, and in Figure 3(b), they have concatenated to two main clusters. Furthermore, when Figures 3(c) and (d) are compared, cluster separability has increased, and all classes occupy the feature space more equally compared to BCL when RCL is applied.
5 Conclusion
The paper proposed RCL for long-tail classification tasks. RCL addresses the balancedness of the learnt feature space, enforces feature compactness and increases the margin width for tail classes. RCL shows a competitive level of accuracy over SOTA methods over three benchmark datasets with imbalance ratios up to 256. RCL was also validated to perform well as a standalone and integrated loss on BCL, an existing supervised contrastive learning loss specifically designed for long-tail classification. It would be an important future work to investigate how much RCL can contribute to optimizing the learnt feature distribution with self-supervised contrastive learning under extreme class imbalance (i.e., higher class imbalance factors than 256).

### v2

@@
 However, when the class imbalance ratio is high, it requires adjusting the supervised contrastive loss to support the tail classes, as the conventional contrastive learning is biased towards head classes by default.
-To this end, we present Rebalanced Contrastive Learning (RCL), an efficient means to increase the long tail classification accuracy by addressing three main aspects: 1.
+To this end, we present Rebalanced Contrastive Learning (RCL), an efficient means to increase the long-tail classification accuracy by addressing three main aspects: 1.
 Feature space balancedness – Equal division of the feature space among all the classes 2.
@@
 We further demonstrate that the performance of RCL as a standalone loss also achieves state-of-the-art level accuracy.
-Keywords:
-Long-tail Classification Contrastive Learning1 Introduction
+[1]organization=The University of Sydney, addressline=Camperdown NSW , city=Sydney, postcode=2050, state=NSW, country=Australia
+1 Introduction
 The data imbalance between different classes usually hinders training an unbiased classifier.
 This is a significant problem, especially when the tail class is of great importance for applications such as pedestrian classification, fraud detection, and spam filtering.
-While approaches such as data resampling and loss reweighting are used to alleviate class imbalance, these approaches, however, increase the accuracy of the tail classes by sacrificing the head class accuracy.
+While approaches such as data resampling and loss reweighting are used to alleviate class imbalance, these approaches increase the tail classes’ accuracy by sacrificing the head class accuracy.
 An important aspect of a robust model is the capability to address the distribution shift from training (long-tail) to unknown test (uniform or inversely long-tailed) distributions.
 Later in some frameworks, cross-entropy loss for long-tail learning is combined with metric learning or contrastive learning.
-Especially work such as [12] demonstrates strong performance improvement when the cross-entropy loss is combined with Supervised Contrastive Learning (SCL).
+According to Jitkrittum et al.
+(2022) strong performance improvement occurs when the cross-entropy loss is combined with Supervised Contrastive Learning (SCL).
 This work demonstrates that learning better embeddings can assist in resolving most of the long-tail classification challenges, as better embeddings can contribute to more generalized classifiers.
-Along this line, one approach is to decouple representation learning and classifier learning [14].
-However, SCL [15] based learnt representation can not eliminate domination by most frequent classes and the decrease of linear separability of the learnt embeddings from head to tail classes unless necessary adjustments for long-tail distributions are conducted.
-Therefore, it is important to improve SCL to handle long-tail data distributions more robustly.
+Along this line, one approach is decoupling representation and classifier learning (Kang et al.
+(2019)).
+However, the representations learned through SCL (Khosla et al.
+(2020)) cannot fully eliminate the dominance of the most frequent classes or maintain the linear separability of embeddings from head-to-tail classes without necessary adjustments for long-tail distributions.
+Therefore, improving SCL to handle long-tail data distributions more robustly is important.
 Zhu et al.
-[29] propose balanced contrastive loss BCL, an improved SCL loss for the context of long-tail learning where the main intuition is to average the attraction and repulsion terms based on the mini-batch class frequency so that each class equally contributes to the optimization process.
+(2022) propose balanced contrastive loss BCL, an improved SCL loss for the context of long-tail learning where the main intuition is to average the attraction and repulsion terms based on the mini-batch class frequency so that each class equally contributes to the optimization process.
 We propose a means to further enhance the quality of the learnt embedding of BCL from the perspectives of balancedness, intra-class compactness, and regularization using Rebalanced Contrastive Learning (RCL).
@@
 Given its simplicity of implementation, it can be incorporated into any SCL method.
-More specifically, we devise the following contributions.
-We propose RCL to optimize the learnt representation of BCL through i) establishing the robustness for distribution shift from train (long tail) to test (uniform) using the adopted class frequency-based SoftMax balancing, ii) enforcing intra-class compactness for tail classes through scalar multiplied features fed to the contrastive loss, and iii) enforcing better class separation through enforced class frequency based embedding margins.
-By conducting experiments on the long-tail versions of the three benchmark datasets: CIFAR10, CIFAR100, and ImageNet, we show that the addition of RCL over BCL improves the top-1 balanced accuracy.
+More specifically, we make the following contributions.
+1.
+We propose RCL to optimize the learnt representation of SCL/BCL through i) establishing the robustness for distribution shift from train (long-tail) to test (uniform) using the adopted class frequency-based SoftMax balancing, ii) enforcing intra-class compactness for tail classes through scalar multiplied features fed to the contrastive loss, and iii) enforcing better class separation through enforced class frequency based embedding margins.
+2.
+By conducting experiments on the long-tail versions of the three benchmark datasets: CIFAR10_Lt, CIFAR100_Lt, and ImageNet_Lt, we show that the addition of RCL over SCL/BCL improves the top-1 balanced accuracy.
 Our ablation study further shows that RCL as a standalone metric also increases the top-1 harmonic mean accuracy, indicating improvements in the least-performing classes.
-We conduct qualitative evaluations of the learnt embeddings and demonstrate that adding RCL to BCL enhances class separability with larger margins through balancedness, intra-class compactness, and regularization.
-2 Related Work
-We explain related work in long-tail classification under three topics: i) Data-level solutions, ii) Algorithm-level solutions, and iii) SCL-based approaches.
-Data-level solutions - Early works in long-tail learning include data-level solutions such as random oversampling, random undersampling [3], square root sampling [19] and progressively balanced sampling [14] that suffer from over-fitting or under-fitting to data.
-Re-weighting methods [8, 25] assign weights on different training samples based on the label or instance (hard or easy samples) and have shown significant accuracy increments over few-shot classes.
+3.
+We conduct qualitative and quantitative evaluations of the learnt embeddings and show that adding RCL to BCL enhances class separability with larger margins through balancedness, intra-class compactness, and regularization.
+2 Related Works
+We explain related work in long-tail classification methods under three topics: i) Data-level solutions, ii) Algorithmic-level solutions, and iii) SCL-based approaches.
+2.1 Data-Level Solutions
+Early works in long-tail learning include data-level solutions such as random oversampling, random undersampling (Chawla et al.
+(2002)), square root sampling (Mahajan et al.
+(2018)) and progressively balanced sampling (Kang et al.
+(2019)) that suffer from over-fitting or under-fitting to data.
+Re-weighting methods (Cui et al.
+(2019); Shu et al.
+(2019)) assign weights on different training samples based on the label or instance (hard or easy samples) and have shown significant accuracy increments over few-shot classes.
 However, re-weighting methods tend to sacrifice the many-shot accuracy.
-Classifier balancing techniques such as fine-tuning the learnt classifier [10], post-hoc calibrations [20, 14], domainshift calibrations [21] alleviate the imbalances of the data.
-However, these methods may not be sufficient under extreme class imbalance ratios and when the number of classes is very high.
-Data augmentation is also used to attain significant accuracy improvements for classes suffering from absolute rarity [22].
-Works such as [5] use generative models to generate new samples for tail classes as a convex combination of existing samples.
-Transferring the information from head classes to tail classes also provides promising results [16].
+Classifier balancing techniques such as fine-tuning the learnt classifier (Guerriero et al.
+(2018)), post-hoc calibrations (Menon et al.
+(2020); Kang et al.
+(2019)), domainshift calibrations (Peng et al.
+(2022)) alleviate the imbalances of the data.
+However, these methods may not be sufficient under extreme class imbalance factor and when the number of classes is very high.
+Data augmentation is also used to improve accuracy for classes suffering from absolute rarity (Perez & Wang (2017)).
+Chu et al.
+(2020) use generative models to generate new samples for tail classes as a convex combination of existing samples.
+Transferring the information from head classes to tail classes also provides promising results (Kim et al.
+(2020)).
 Yet, these methods can not completely eliminate the tail class over-fitting and the dominance of head classes.
-Algorithmic-level solutions - Under algorithmic level solutions integrating a per-class margin [20, 23] into the cross-entropy loss is a popular direction of work.
+2.2 Algorithmic-Level Solutions
+Under algorithmic level solutions integrating a per-class margin (Menon et al.
+(2020); Ren et al.
+(2020)) into the cross-entropy loss is a popular work direction.
 In most cases, the logit margin is set as inversely proportional to the class frequency to enforce a larger margin for tail classes to enhance generalization.
@@
 Therefore, it is important to use techniques to optimize the learned feature distribution to avoid the diffusion of the rare classes.
-Representation optimization with SCL - The contrastive learning-based methods have been used recently in the context of long-tail learning as they contribute to better feature representation by aligning the same class pairs together and repulsing the different class pairs.
-In [18], authors propose a novel technique to improve SCL so that it computes optimal class centres to balance the features space evenly.
-Then, they associate the generated class centres with actual class IDs considering mutual closeness.
+2.3 Representation Optimization With SCL
+The contrastive learning-based methods have been used recently in the context of long-tail learning as they contribute to better feature representation by aligning the same class pairs together and repulsing the different class pairs.
+Li et al.
+(2022) propose a novel technique to improve SCL to compute optimal class centres and evenly balance the features space.
+Then, they associate the generated class centres with actual class IDs, considering mutual closeness.
 However, the results here could be suboptimal since there is no deterministic strategy to associate the class centres with the true classes.
 Song et al.
-[26] propose an improved SCL framework denoted as prototypical supervised contrastive (PSC) learning where it learns prototypes for each class; subsequently, each instance is to be pushed towards the prototypes of its corresponding class and pulled away from the other class prototypes.
-In this approach, tail class instances may have a larger effect on the model optimization compared to head class instances, sacrificing the head class accuracy.
-To address this issue, [11] proposes subclass-balancing contrastive learning where the loss is defined on subclasses equal to the tail class size inside the head classes through clustering.
-Here, the authors address issues with balancing the SCL loss minimized along with the CE loss in the context of long-tail classification scenarios with higher class imbalance.
-They focus on increasing the balancedness, intra-class compactness, and tail-class margins through rebalancing the BCL loss.
+(2022) introduce an improved SCL framework denoted as prototypical supervised contrastive (PSC) learning where it learns prototypes for each class; subsequently, each instance is to be pushed towards the prototypes of its corresponding class and pulled away from the other class prototypes.
+In this approach, tail class instances may have a larger effect on the model optimization than head class instances, sacrificing the head class accuracy.
+To address this issue, Hou et al.
+(2023) showcase subclass-balancing contrastive learning where the loss is defined on subclasses equal to the tail class size inside the head classes through clustering.
+Here, the authors balance the SCL loss accordingly to support long-tail classification scenarios with higher class imbalance and subsequently balanced SCL loss is jointly minimized with CE loss.
+Our work focus on a novel yet simpler approach based on BalanceSoftmax concept to efficiently balance the SCL component so that it can significantly enhance the balancedness, intra-class compactness, and tail-class margins through rebalancing the BCL loss.
 3 Methodology
@@
 Learned embedding corresponding to the input is .
-Further, we denote the batch frequency of class y by and the learnt class prototype by as defined in [29].
+Further, we denote the batch frequency of class y by and the learnt class prototype by as defined in the work of Zhu et al.
+(2022).
 3.2 Preliminaries
-CE Based Classification: Our work makes use of fundamental Cross Entropy (CE) modification methods for long-tail learning.
-Firstly, as the SoftMax function is inherently biased under long-tail learning, a Balanced SoftMax function [23] is proposed to address the train-to-test distribution shift from a probabilistic perspective.
+3.2.1 CE Based Classification:
+Our work makes use of fundamental Cross Entropy (CE) modification methods for long-tail learning.
+Firstly, as the SoftMax function is inherently biased under long-tail learning, a Balanced SoftMax function (Ren et al.
+(2020)) is proposed to address the train-to-test distribution shift from a probabilistic perspective.
 The modified loss is denoted in Equation 1.
@@
 The second interesting approach is logit adjustment (LA) proposed by Menon et al.
-[20] that allows larger margins for tail classes, which can be denoted with Equation 2.
+(2020) that allows larger margins for tail classes, which can be denoted with Equation 2.
 Here denotes the probability of class y.
 This model is more favourable as provides the degree of freedom to adjust for different datasets and class imbalance levels.
-SCL Integration: Above LA cross entropy classification module combined with a supervised contrastive learning module to enhance the long-tail classification accuracy.
-The supervised contrastive learning component is balanced through class averaging and class complements [29].
+3.2.2 SCL Integration:
+Above LA cross entropy classification module combined with a supervised contrastive learning module to enhance the long-tail classification accuracy.
+The supervised contrastive learning component is balanced through class averaging and class complements (Zhu et al.
+(2022)).
 The new loss is denoted as balanced contrastive loss BCL.
@@
 3.3 Feature Representation Optimization With RCL
-Our RCL-based approach addresses three key aspects in the embedding space that can directly contribute to increasing classification accuracy: 1) Balanced feature space, 2) Intra-class compactness, and 3) Assigning larger margins to ambiguous classes.
+Our RCL-based approach addresses three key aspects in the embedding space that can directly contribute to increasing accuracy: 1) Balanced feature space, 2) Intra-class compactness, and 3) Assigning larger margins to ambiguous classes.
 These key aspects are expected to provide more support for tail class classification.
@@
 In Figure 2, we illustrate the pipeline for the RCL-based learning.
-In this work, we apply the RCL loss to the BCL loss.
-However, RCL loss can be integrated with any given similar SCL pipeline.
+In this work, we analyse the RCL directly applied to the SCL loss, and secondly, RCL is applied to the BCL loss.
 3.3.1 Balancing Feature Space
-It was noted that unsupervised contrastive learning methods are more robust to highly imbalanced class distributions compared to supervised contrastive learning [13].
+It was noted that unsupervised contrastive learning methods are more robust to highly imbalanced class distributions compared to supervised contrastive learning (Kang et al.
+(2020)).
 The reason for this phenomenon is that the unsupervised CL methods generate a balanced feature space for all the classes compared to SCL.
-To this end, some research focuses on pushing the learnt embeddings into pre-defined class centres to obtain the balance [18].
+To this end, some research focuses on pushing the learnt embeddings into pre-defined class centres to obtain the balance (Li et al.
+(2022)).
 In this work, we try to obtain balancedness by balancing the SCL loss sufficiently.
@@
 Where is the frequency of the class y in the complete training set.
-The RCL modification can also be implemented on BCL to enhance the balancedness as in Equation 5.
-The fraction denoted in red is the class averaged similarity scores.
 3.3.2 Enforcing Feature Compactness
 Minimizing the intra-class distance among the same class embeddings is a main aspect of interest as it reduces the deviation of the features across the decision boundary and enhances class discriminativeness.
-Feature Clusters Compression (FCC) [17] is applied previously to decrease the diffusion of backbone features by compressing learnt embeddings.
-This is simply achieved by multiplying original learnt features by a selected scaling factor and using the multiplied features in computing the cross entropy loss for classifier learning.
+Feature Clusters Compression (FCC) (Li et al.
+(2023)) is applied previously to decrease the diffusion of backbone features by compressing learnt embeddings.
+This is achieved by multiplying original learnt features by a selected scaling factor and using the multiplied features in computing the cross entropy loss for classifier learning.
 In this work, we modify this approach and integrate the feature compression component directly into the RCL instead of sending it through the classifier as in CE minimization, i.e., .
@@
 The scaling factor and value is selected based on the class id.
-At the The validation accuracy is obtained, and if the validation per class accuracy is lower than , a novel scaling factor is applied for the corresponding classes for the rest of the classes, the factor is set to 1 (only two different values are used for the entire data set).
+At the , the validation accuracy is obtained, and if the validation per class accuracy is lower than , a novel scaling factor is applied for the corresponding classes and for the rest of the classes, the factor is set to 1 (only two values are used for the entire dataset).
 The scaling factor used on the underperforming classes is selected based on a grid search.
+The RCL modification with feature compression can be implemented on BCL to enhance the balancedness as in Equation 5.
+The fraction in red is the class averaged similarity scores.
 3.3.3 Regularization
@@
 Therefore, the data-dependent properties like class margins can be useful.
-Enforcing a larger margin for a class can be interpreted as regularization [2] because the inverse of the minimum margin among all the examples is linked with the standard generalization error bounds.
-According to [12], enforcing logit margins is insufficient under extreme imbalance scenarios, and therefore, embedding margins are also used to enhance the regularization of the minority classes.
-We demonstrate the importance of the RCL in the context of enforced embedding margins by reformulating it as in Equation 6 followed by Equation 7.
+Enforcing a larger margin for a class can be interpreted as regularization (Cao et al.
+(2019)) because the inverse of the minimum margin among all the examples is linked with the standard generalization error bounds.
+Jitkrittum et al.
+(2022) demonstrate that enforcing logit margins is insufficient under extreme imbalance scenarios, and therefore, embedding margins are required to enhance the regularization of the minority classes.
+We demonstrate the RCL’s importance in enforced embedding margins by reformulating it as in Equation 6 followed by Equation 7.
 According to this format, it is evident that margin is implicitly enforced for the similarity score between the same class and dissimilar class embeddings based on the class frequency.
@@
 Figure 1(b) denotes a 2D classification problem which portrays the ideally expected distribution of the learnt features under an imbalanced data distribution.
-All the classes occupy a similar area in the feature space, and the features are well compressed towards the class centre while providing larger margins for the rare classes.
-4 Experiments
-| LC | SCL | BCL | RCL | Top-1 Accuracy (Arithmetic) | Top-1 Accuracy (Harmonic) |
-[…]
-[…]
-[…]
-[…]
-[…]
-| Dataset | CIFAR10 | CIFAR100 | ImageNet |
-| Imbalance Ratio | 100 | 100 | 256 |
-| CE | 70.4 | 38.5 | 41.6 |
-| LDAM+DRW [2] | 77.2 | 42.2 | 49.8 |
-| LogitAdjust [20] | 80.2 | 43.9 | 50.9 |
-| CB+Focal [8] | 74.6 | 39.9 | - |
-| BALMS [23] | 84.9 | 50.8 | - |
-| Two Stage Methods | |||
-| Decoupled [14] | 78.5 | 41.3 | 50.6 |
-| DRO_LT [24] | 80.5 | 47.1 | 53.5 |
-| Ensemble Methods | |||
-| RIDE [28] | - | 49.1 | 55.4 |
-| ACE [1] | 81.4 | 49.6 | 54.7 |
-| CE+ other Loss | |||
-| ELM [12] | 78.0 | 45.8 | 50.6 |
-| CE+ Supervised Contrastive Loss | |||
-| Hybrid_spc [27] | 81.4 | 46.7 | - |
-| BCL | 84.3 | 51.9 | 56.2 |
-| RCL (ours) | 86.0 | 52.2 | 55.4 |
-| BCL+RCL (ours) | 86.2 | 52.1 | 56.3 |
-To validate our findings on representation learning optimization with RCL, we use three benchmark datasets containing significant class imbalance, exhibiting long-tail distribution and involving a large number of classes.
-We next describe the datasets, experiment design, and performance metrics, followed by comparisons of the results with the SOTA techniques and ablation studies.
+All the classes occupy a similar area in the feature space, and the features are well compressed towards the class centre with larger margins for the rare classes.
+4 Experimental Setup and Results
+| LC | SCL | BCL | RCL | Arithmetic | Harmonic |
+| ✓ | - | - | - | 44.0 | 35.2 |
+| ✓ | ✓ | - | - | 51.1 | 39.2 |
+| ✓ | ✓ | ✓ | - | 51.9 | 40.2 |
+| ✓ | ✓ | - | ✓ | 52.2 | 42.4 |
+| ✓ | ✓ | ✓ | ✓ | 52.1 | 41.1 |
+| LC | SCL | BCL | RCL | Arithmetic | Harmonic |
+| ✓ | - | - | - | 81.2 | 80.5 |
+| ✓ | ✓ | - | - | 83.5 | 83.1 |
+| ✓ | ✓ | ✓ | - | 84.3 | 83.7 |
+| ✓ | ✓ | - | ✓ | 86.0 | 85.3 |
+| ✓ | ✓ | ✓ | ✓ | 86.2 | 85.7 |
+| Dataset | CIFAR10_Lt | CIFAR100_Lt | ||||
+| Imbalance Ratio | 100 | 50 | 10 | 100 | 50 | 10 |
+| Cross Entropy (CE) | 70.4 | 74.8 | 86.4 | 38.5 | 43.9 | 55.7 |
+| LDAM+DRW (Cao et al. (2019)) | 77.2 | 81.0 | 88.2 | 42.2 | 46.6 | 58.7 |
+| LogitAdjust (Menon et al. (2020)) | 80.2 | - | - | 43.9 | - | 57.4 |
+| CB+Focal (Cui et al. (2019)) | 74.6 | 79.4 | 87.1 | 39.9 | 45.2 | 58.0 |
+| BALMS (Ren et al. (2020)) | 84.9 | - | 91.3 | 50.8 | - | 63.0 |
+| Two Stage Methods | ||||||
+| Decoupled (Kang et al. (2019)) | 78.5 | - | 91.1 | 41.3 | - | 63.4 |
+| DRO_LT (Samuel & Chechik (2021)) | 80.5 | - | - | 47.1 | 57.6 | 63.4 |
+| Ensemble Methods | ||||||
+| RIDE (Wang et al. (2020)) | - | - | - | 49.1 | - | 61.8 |
+| ACE (Cai et al. (2021)) | 81.4 | 84.9 | - | 49.6 | 51.9 | - |
+| CE+ Other Losses | ||||||
+| BBN (Zhou et al. (2020)) | 79.8 | 82.2 | - | 42.6 | 47.0 | - |
+| CE+Supervised Contrastive Loss | ||||||
+| Hybrid_spc (Wang et al. (2021)) | 81.4 | 83.7 | 91.1 | 46.7 | 48.9 | 63.1 |
+| TSC (Li et al. (2022)) | 79.7 | 82.9 | 88.7 | 43.8 | 47.4 | 59.0 |
+| BCL | 84.3 | 87.2 | 91.1 | 51.9 | 56.6 | 64.9 |
+| RCL (ours) | 86.0 | 88.5 | 91.7 | 52.2 | 56.8 | 65.6 |
+| BCL+RCL (ours) | 86.2 | 88.4 | 92.0 | 52.1 | 56.0 | 65.4 |
+| Dataset | ResNet50 | ResNext50 |
+| CE | 41.6 | 44.4 |
+| LDAM+DR | 49.8 | - |
+| LogitAdjust | 50.9 | - |
+| CB+Focal | 46.8 | 45.3 |
+| BalSoftMax | - | 51.4 |
+| Two Stage Methods | ||
+| Decoupled | 50.6 | 49.6 |
+| DRO_LT | 53.5 | 53.5 |
+| Ensemble Methods | ||
+| RIDE | 55.4 | 56.8 |
+| ACE | 54.7 | 56.6 |
+| CE+Other Losses | ||
+| ELM | 50.6 | - |
+| CE+Supervised Contrastive Loss | ||
+| BBN | 48.3 | 49.3 |
+| BCL | 56.2 | 57.4 |
+| BCL+RCL (ours) | 56.3 | 57.5 |
+To validate our findings on representation learning optimization with RCL, we use three benchmark datasets with significant class imbalance, exhibiting long-tail distribution and involving a large number of classes.
+Next, we describe the datasets, experiment design followed by comparisons of the results with the SOTA and ablation study.
 4.1 Datasets
-We use the long-tail versions of the benchmark datasets: CIFAR10, CIFAR100, and ImageNet.
+We use the long-tail versions of the benchmark datasets: CIFAR10_Lt, CIFAR100_Lt, and ImageNet_Lt.
 The level of class imbalance is measured by the Class Imbalance Factor (IF), which is the ratio between the most frequent class frequency and the least frequent class frequency.
 We set the maximum possible imbalance factors for the datasets, leaving at least 5-6 instances per tail class.
-Doing so, we obtain the following experiment settings; i) CIFAR10: No.
-of classes 10, IF - 100, majority class frequency - 5,000, ii) CIFAR100: No.
-of classes 100, IF - 100, majority class frequency - 500, iii) ImageNet: No.
+Doing so, we obtain the following experiment settings; i) CIFAR10_Lt: No.
+of classes 10, IF - 100, majority class frequency - 5,000, ii) CIFAR100_Lt: No.
+of classes 100, IF - 100, majority class frequency - 500, iii) ImageNet_Lt: No.
 of classes 1,000, IF - 256, majority class frequency - 1,280.
@@
 Models are trained with a batch size of 256.
-Learning rate for CIFAR10 and 100 are set to 0.2 and for ImageNet 0.1.
-Similar to [7] and [29], we use Autoaugment [6] and Cutout [9] for augmenting the features fed to the classifier.
-SimAugment [4] is used for the features fed to the contrastive loss.
-The model training is conducted on the Amazon Cloud with RONIN mediation service with 8xV100 GPUs.
+Learning rate for CIFAR10_Lt and CIFAR100_Lt are set to 0.2 and for ImageNet_Lt 0.1.
+Similar to work of Cui et al.
+(2021) and Zhu et al.
+(2022), we use Autoaugment (Cubuk et al.
+(2019)) and Cutout (DeVries & Taylor (2017)) for augmenting the features fed to the classifier.
+SimAugment (Chen et al.
+(2020)) is used for the features fed to the contrastive loss.
+The model was trained on a Amazon Cloud VM with 8xV100 GPUs.
 4.3 Ablation Study
-We first conduct an ablation study on CIFAR100 with IF set to 100 to analyse the performance benefits of RCL.
-The overall balanced accuracy (averaged per class) is reported in Table 1.
-As small values heavily influence arithmetic means in obtaining the overall class accuracy, we also report the harmonic mean to reflect the performance of the worst-performing classes.
-Initially, the dataset is tested for cross-entropy loss with LA [20].
-Where the logit adjustment parameter is set to 1.3 through grid search.
+We conducted an ablation study on CIFAR100_Lt and CIFAR10_Lt with IF set to 100 to analyse the performance benefits of RCL.
+The overall balanced accuracy (averaged per class) is reported in Table 1 and Table 2, respectively.
+As large values heavily influence arithmetic means in obtaining the overall class accuracy, we also report the harmonic mean to reflect the performance of the worst-performing classes.
+Initially, the dataset is tested for cross-entropy loss with LA (Menon et al.
+(2020)).
+Where the logit adjustment parameter is set to 1.3 for RCL through grid search.
 Then, the accuracy is evaluated by combining standard supervised contrastive loss with setting to 2 and to 1 and using SimAugment for augmentation.
-The compression factor is set to 0.005 for the classes with validation accuracy less than , and for the other classes, the factor is set to 1.
+The compression factor is set to 1 for the classes with validation accuracy less than ; for the other classes, the factor is set to 0.005.
 Compression is applied from the epoch.
-Results in Table 1 show adding RCL to standard SCL would increase top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4).
+CIFAR100_Lt results in Table 1 show adding RCL to standard SCL would increase top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4).
 When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5).
 SCL+RCL demonstrates the highest harmonic mean accuracy of all the combinations.
+Similarly, Table 2 shows the results for CIFAR10_Lt where adding RCL to SCL increases top-1 accuracy by and harmonic mean accuracy by (rows 2 and 4).
+When RCL is added to BCL, the top-1 accuracy increases by and harmonic mean accuracy by (rows 3 and 5).
+SCL+BCL+RCL demonstrates the highest harmonic mean accuracy of all.
 4.4 Performance Analysis
 4.4.1 Baselines
-We compare our proposed framework with twelve SOTA models.
+We compare our proposed framework with 13 SOTA models under five main categories.
 Balanced overall test accuracy (arithmetic mean over per-class accuracy) is used for comparison.
-The selected baselines are: 1) standard cross entropy minimisation, 2) LDAM-DRW where logit margin is enforced depending on the class frequency and reweighting scheme is used in optimization [2], 3) Class balanced focal loss [8] which assigns a frequency based weighting on per class losses, 4) Additive logit margin based on the class frequency [20], 5) BALMS - Balanced SoftMax function Equation [23] to address the distribution shift integrated with a meta sampler that can compute the optimal class sample rate, 6) Decoupled learning which enhances the classification accuracy through fine-tuning a learnt classifier with class-balanced sampling or classifier weight normalization [14], 7) robustness theory based loss to enhance the classification and feature representation [24], 8) RIDE [28] and 9) ACE [1] which are multi expert ensemble models, 10) ELM [12] which enforces logit and embedding margins, and 11) Hybrid_spc [27] prototypical supervised contrastive (PSC) learning.
+The selected categories of algorithmic level solutions are 1) Standard and modified (logit adjusted) CE minimization, 2) Two-stage learning methods where representation learning and classifier fine-tuning are handled in different stages; 3) Ensemble methods where multiple experts combined to enhance the head and tail class accuracy simultaneously, 4) CE loss combined with an additional loss to handle representation learning explicitly, 5) CE loss combined with SCL loss to enhance representation learning.
 4.4.2 Results
 We show the results in Table 4.
-As can be seen, standalone RCL demonstrates , improvement over CIFAR10 and 100, respectively, compared to BCL, which is the next best model.
-Furthermore, BCL+RCL has consistent improvements , , and over all the datasets in comparison to BCL.
-In summary, standalone RCL and BCL+RCL outperform all the other baselines apart from one scenario where BCL performs best (i.e., ImageNet).
+As can be seen, standalone RCL demonstrates satisfactory performance compared to the SOTA methods.
+The improvement over BCL is 1.7%,1.3% and 0.6% for CIFAR10_Lt for IF = 100, 50 and 10, respectively.
+Similar to CIFAR100_Lt, the values are 0.3%,0.2% and 0.7%.
+Even though the CIFAR100_Lt improvement over BCL is small compared to the other popular SCL methods, RCL has a competitive performance.
+Furthermore, BCL+RCL has consistent improvements 1.9%, 1.2%, and 0.9% for CIFAR10_Lt.
+And for CIFAR100_Lt, the corresponding improvements are 0.2% for IF100 and 0.5 % IF=10.
+In summary, BCL performs better for CIFAR100_Lt IF=50.
+For Imagenet_Lt, BCL+RCL outperforms BCL by 0.1% in both backbones according to Table 4.
+| Dataset (IF=100) | BCL | BCL+RCL | ||
+| CHI | DBI | CHI | DBI | |
+| CIFAR100_Lt | 3,962.9 | 0.97 | 4,091.8 | 0.94 |
+| CIFAR10_Lt | 84.1 | 4.1 | 105.8 | 4.0 |
 4.4.3 Embedding Space Analysis
-In Figure 3, we show the TSNE plots of the learnt representation of the model for CIFAR10 and ImageNet.
-BCL+RCL demonstrates a better class separation and balanced division of the feature space compared to BCL.
-As seen from Figure 3(a), in BCL, tail class embeddings are grouped into sub-clusters and lie far apart.
-Once the features compression is added, all embeddings corresponding to the same tail class are grouped together, providing better intra-class compactness and balancedness in the space in Figure 3(b).
-For example, in Figure 3(a), tail classes 8 and 9, represented by orange and cyan, have two sub-clusters each, and in Figure 3(b), they have concatenated to two main clusters.
-Furthermore, when Figures 3(c) and (d) are compared, cluster separability has increased, and all classes occupy the feature space more equally compared to BCL when RCL is applied.
+According to the quantitative analysis of the RCL method performance, it is evident that standalone RCL applied to the standard SCL demonstrates SOTA competitive accuracy levels.
+Apart from this, RCL demonstrates an interesting aspect when applied to the BCL.
+Even though the accuracy improvement over BCL is relatively small, adding the RCL to the BCL enhances the learnt feature distribution.
+In Figure 3, we show the TSNE plots of the learnt representation of the model for CIFAR10_Lt and CIFAR100_Lt.
+BCL+RCL demonstrates a better class separation and balanced feature space division than BCL.
+As seen from Figure 3(a), in BCL, tail class embeddings are grouped into sub-clusters and are far apart.
+Once the features compression is added, all embeddings corresponding to the same tail class are grouped together, providing better intra-class compactness and balancedness in the space.
+Figure 3(b) visually showcase the better class separation and balanced space distribution for all classes except for 1 and 9 compared to Figure 3(a); furthermore, when Figures 3(c) and (d) are compared, cluster balancedness has increased as a whole, as all classes occupy the feature space more equally compared to BCL when RCL is applied.
+In addition, compactness for classes 40 and 90 appears to increase significantly.
+To further validate this phenomenon, we compute the Calinski–Harabasz index (Caliński & Harabasz (1974)) for the learnt features of the validation set.
+A higher index value indicates better cluster cohesion and cluster separation.
+Table 5 indicates the index values for CIFAR10_Lt and CIFAR100_Lt.
+Where we can see a consistent increment in the Calinski–Harabasz index for both datasets for BCL+RCL over BCL.
+Davies-Bouldin Index assesses the clustering quality where a lower value indicates more compact and well-separated clusters.
+According to the table, we can see that both datasets have a consistently lower value for BCL+RCL over BCL, which implies better cluster quality.
+The index is relatively insensitive to the choice of distance metric.
+Based on this analysis, we can conclude that RCL has the capability to optimize the distribution of the learnt feature simultaneously with the classification accuracy.
+The feature distribution is improved in balancedness, class compactness (cohesion) and regularization (larger margins, better class separation).
 5 Conclusion
@@
 RCL addresses the balancedness of the learnt feature space, enforces feature compactness and increases the margin width for tail classes.
-RCL shows a competitive level of accuracy over SOTA methods over three benchmark datasets with imbalance ratios up to 256.
+RCL shows a competitive level of accuracy over SOTA methods over three benchmark datasets with imbalance factors up to 256.
 RCL was also validated to perform well as a standalone and integrated loss on BCL, an existing supervised contrastive learning loss specifically designed for long-tail classification.
-It would be an important future work to investigate how much RCL can contribute to optimizing the learnt feature distribution with self-supervised contrastive learning under extreme class imbalance (i.e., higher class imbalance factors than 256).
+It would be an important future work to investigate how much RCL can contribute to optimizing the learnt feature distribution with self-supervised contrastive learning under extreme class imbalance (i.e., IF than 256).
+Acknowledgments
+This research was conducted by the University of Sydney for National Intelligence Postdoctoral Grant (project number NIPG-2022-006) and funded by the Australian Government.